### v1

@@ 2 $S$-Primary Decomposition in $S$-Noetherian Ring
(Existence of $S$-Primary Decomposition) Let $R$ be an $S$-Noetherian ring. Then every proper ideal (disjoint with $S$) of $R$ can be written as a finite intersection of $S$-primary ideals.
@@
Let $E$ be a collection of ideals of $R$ which can not be written as a finite intersection of $S$-irreducible ideals. We wish to show $E=\emptyset$. On contrary suppose $E\neq\emptyset$. Since $R$ is an $S$-Noetherian ring, by [9, Theorem 2.3], there exists an $S$-maximal element in $E$, say $I$. Evidently, $I$ is not an $S$-irreducible ideal, and so $I$ is not an irreducible ideal. This implies that $I=J\cap K$ for some ideals $J$ and $K$ of $R$ with $I\neq J$ and $I\neq K$. Since $I$ is not $S$-irreducible, $sJ\nsubseteq I$ and $sK\nsubseteq I$ for all $s\in S$. Now we claim that $J$, $K\notin E$. For this, suppose that $J$ and $K$ are in $E$, then there exist $s^{\prime},s^{\prime\prime}\in S$ such that $s^{\prime}J\subseteq I$ and $s^{\prime\prime}K\subseteq I$ and $I\subset J$, $I\subset K$ which is not possible. Therefore $J,K\notin E$. This implies that $J$ and $K$ can be written as a finite intersections of $S$-irreducible ideals. Consequently, $I$ can also be written as a finite intersection of $S$-irreducible ideals since $I=J\cap K$, a contradiction as $I\in E$. Thus $E=\emptyset$, i.e., every proper ideal of $R$ can be written as finite intersection of $S$-irreducible ideals. But by Theorem 10, an $S$-irreducible ideal is $S$-primary. Therefore every ideal of $R$ can be written as a finite intersection of $S$-primary ideals. ∎
Recall that if $I$ is any ideal of $R$, the radical of $I$ is $rad(I)=\{x\in R\hskip 2.84544pt|\hskip 2.84544ptx^{n}\in I\hskip 2.84544ptfor\hskip 2.84544ptsome\hskip 2.84544ptn\geq 0\}$. If $I=rad(I)$, then $I$ is called radical ideal.
As an application of this theorem, we provide the following result.
@@
Let $R$ be an $S$-Noetherian ring. Then every radical ideal $I$ (disjoint with $S$) of $R$ is the intersection of finitely many $S$-prime ideals.
@@
By Theorem 11, there exist finitely many $S$-primary ideals $Q_{1},Q_{2},\ldots,Q_{r}$ such that $I=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$. Also by [14, Proposition 2.5], $rad({Q_{i}})=P_{i}$ is $S$-prime for $i=1,2,\ldots,r$. Consequently, $I=rad(I)=P_{1}\cap P_{2}\cap\cdots\cap P_{r}$. This completes the proof. ∎
@@
Let $R$ be an $S$-Noetherian ring and $I$ be an ideal of $R$. Then by Theorem 11, $I$ has $S$-primary decomposition. Suppose $I=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$ is an $S$-primary decomposition of $I$. Since each $Q_{i}$ is $S$-primary, so by [16, Proposition 2.6], each $S^{-1}Q_{i}$ is primary. Consequently, $S^{-1}I=S^{-1}Q_{1}\cap S^{-1}Q_{2}\cap\cdots\cap S^{-1}Q_{r}$ is a primary decomposition of $S^{-1}I$
To prove $S$-version of the $1st$ uniqueness theorem for $S$-Noetherian rings, we need the following lemmas:
Lemma 14.
@@
On contrary suppose $sQ_{k}\nsubseteq P$ for all $s\in S$ and $k=1,2,\dots,n$. Then for each $k$, there exists $x_{k}\in Q_{k}$ such that $sx_{k}\notin P$ for all $s\in S$. Evidently, $x_{1}x_{2}\ldots x_{n}\in Q_{1}Q_{2}\ldots Q_{n}$. This implies that $(sx_{1})(sx_{2})\cdots(sx_{n})\in Q_{1}Q_{2}\ldots Q_{n}\subseteq\bigcap\limits_{k=1}^{n}Q_{k}\subseteq P$. Then there exist $s^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $s^{\prime}sx_{j}\in P$ since $P$ is an $S$-prime ideal, a contradiction. Hence there exist $s\in S$ and $k\in\{1,2,\ldots,n\}$ such that $sQ_{k}\subseteq P$. Finally, if $P=\bigcap\limits_{k=1}^{n}Q_{k}$, then by above argument there exist $s\in S$ and $k\in\{1,2,\ldots,n\}$ such that $sQ_{k}\subseteq P\subseteq Q_{k}$. ∎
@@
Let $Q$ be an $S$-primary ideal and $x$ an element of $R$ such that $sx\notin Q$ for all $s\in S$. Then $(Q:sx)$ is $S$-primary, and there exists $t\in S$ such that $t(rad(Q:sx))\subseteq rad(Q)\subseteq rad(Q:sx)$ for all $s\in S$.
@@
Suppose $(Q:sx)\cap S\neq\emptyset$ for some $s\in S$. Then there exists $t\in S$ such that $t\in(Q:sx)$ which imples that $tsx\in Q$, a contradiction. Thus $(Q:sx)\cap S=\emptyset$ for all $s\in S.$ Now let $s\in S$ and $yz\in(Q:sx)$ be such that $s^{\prime}y\notin rad(Q:sx)$ for all $s^{\prime}\in S$. Then $sxyz\in Q$. Since $Q$ is $S$-primary and $s^{\prime}y\notin rad(Q:sx)$ for all $s^{\prime}\in S$, there exists $s^{\prime\prime}\in S$ such that $s^{\prime\prime}sxz\in Q$. Consequently, $s^{\prime\prime}z\in(Q:sx)$. Thus $(Q:sx)$ is $S$-primary. Further, let $w\in rad(Q:sx)$. Then $w^{n}\in(Q:sx)$ for some $n\in\mathbb{N}$. This implies that $sxw^{n}\in Q$. Since $Q$ is an $S$-primary, there exists $t\in S$ such that $(tw^{n})^{k}\in Q$ for some $k$ as $tsx\notin Q$. This concludes that $(tw)^{nk}=t^{nk-k}(tw^{n})^{k}\in Q$ which implies that $tw\in rad(Q)$. Thus $t(rad(Q:sx))\subseteq rad(Q)$. Also $Q\subseteq(Q:sx)$ implies that $rad(Q)\subseteq rad(Q:sx)$, and hence $t(rad(Q:sx))\subseteq rad(Q)\subseteq rad(Q:sx)$. ∎
@@
[…]

### v2

@@ 2 $S$-Primary Decomposition in $S$-Noetherian Ring
(Existence of $S$-Primary Decomposition) Let $R$ be an $S$-Noetherian ring. Then every proper ideal of $R$ disjoint from $S$ can be written as a finite intersection of $S$-primary ideals.
@@
Let $E$ be the collection of ideals of $R$ which are disjoint from $S$ and can not be written as a finite intersection of $S$-primary ideals. We wish to show $E=\emptyset$. On the contrary suppose $E\neq\emptyset$. Since $R$ is an $S$-Noetherian ring, by [8, Theorem 2.3], there exists an $S$-maximal element in $E$, say $I$. Evidently, $I$ is not an $S$-primary ideal, by Theorem 8, $I$ is not an $S$-irreducible ideal, and so $I$ is not an irreducible ideal. This implies that $I=J\cap K$ for some ideals $J$ and $K$ of $R$ with $I\neq J$ and $I\neq K$. Since $I$ is not $S$-irreducible, $sJ\nsubseteq I$ and $sK\nsubseteq I$ for all $s\in S$. Now we claim that $J$, $K\notin E$. For this, if $J$ (respectively, $K$) belongs to $E$, then since $I$ is an $S$-maximal element of $E$ and $I\subset J$ (respectively, $I\subset K$), there exists $s^{\prime}$ (respectively, $s^{\prime\prime}$) from $S$ such that $s^{\prime}J\subseteq I$ (respectively, $s^{\prime\prime}K\subseteq I$). This is not possible, as $I$ is not $S$-irreducible. Therefore $J,K\notin E$. This implies that $J$ and $K$ can be written as a finite intersection of $S$-primary ideals. Consequently, $I$ can also be written as a finite intersection of $S$-primary ideals since $I=J\cap K$, a contradiction as $I\in E$. Thus $E=\emptyset$, i.e., every proper ideal of $R$ disjoint from $S$ can be written as a finite intersection of $S$-primary ideals. ∎
Recall that if $I$ is any ideal of $R$, the radical of $I$ is $rad(I)=\{x\in R\hskip 2.84544pt|\hskip 2.84544ptx^{n}\in I\hskip 2.84544ptfor\hskip 2.84544ptsome\hskip 2.84544ptn\textgreater 0\}$. If $I=rad(I)$, then $I$ is called a radical ideal.
@@
Let $R$ be an $S$-Noetherian ring. Then every radical ideal $I$ (disjoint from $S$) of $R$ is the intersection of finitely many $S$-prime ideals.
@@
By Theorem 9, there exist finitely many $S$-primary ideals $Q_{1},Q_{2},\ldots,Q_{r}$ such that $I=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$. Also by [13, Proposition 2.5], $rad({Q_{i}})=P_{i}$ is $S$-prime for $i=1,2,\ldots,r$. Consequently, $I=rad(I)=P_{1}\cap P_{2}\cap\cdots\cap P_{r}$. This completes the proof. ∎
@@
Let $R$ be a ring, and let $S$ be a multiplicative set of $R$. Let $I$ be an ideal of $R$ such that $I\cap S=\emptyset$. Suppose that $I$ admits an $S$-primary decomposition. (In Theorem 9, we have shown that if $R$ is $S$-Noetherian, then any ideal $I$ of $R$ disjoint from $S$ admits an $S$-primary decomposition). Let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be an $S$-primary decomposition of $I$ with $Q_{i}$ is a $P_{i}$-$S$-primary ideal of $R$ for each $i\in\{1,2,\ldots,n\}$. Then, by [13, Proposition 2.7], $S^{-1}Q_{i}$ is $S^{-1}P_{i}$-primary, $S(I)=\bigcap\limits_{i=1}^{n}S(Q_{i})$ is a primary decomposition with $S(Q_{i})$ is a $S(P_{i})$-primary for each $i\in\{1,\ldots,n\}$. Let $1\leq i\leq n$. Let $s_{i}\in S$ be such that it satisfies the $S$-primary property of $Q_{i}$. Notice that $(P_{i}:s_{i})$ is a prime ideal of $R$ and $(Q_{i}:s_{i})$ is a $(P_{i}:s_{i})$-primary ideal of $R$. Let $s=\prod_{i=1}^{n}s_{i}$. Then $s\in S$. Observe that $(Q_{i}:s_{i})=(Q_{i}:s)$, $(P_{i}:s_{i})=(P_{i}:s)$, $S(Q_{i})=(Q_{i}:s)$, and $S(P_{i})=(P_{i}:s)$. From $I=\bigcap\limits_{i=1}^{n}Q_{i}$, it follows that $(I:s)=\bigcap\limits_{i=1}^{n}(Q_{i}:s)=\bigcap\limits_{i=1}^{n}S(Q_{i})=S(I)$. Let $k$ of the $S(P_{1}),\ldots,S(P_{n})$ be distinct. After a suitable rearrangement of $\{1,\ldots,n\}$, we can assume without loss of generality that $S(P_{1}),\ldots,S(P_{k})$ are distinct among $S(P_{1}),\ldots,S(P_{n})$. Let $A_{1}=\{j\in\{1,\ldots,n\}\mid S(Q_{j})~is~S(P_{1})$-primary}\}, …, $A_{k}=\{j\in\{1,\ldots,n\}\mid S(Q_{j})~is~S(P_{k})$-primary }\}.
From the above discussion, it is evident that $1\in A_{1},\ldots,k\in A_{k}$, and $\{1,\ldots,n\}=\bigcup\limits_{t=1}^{k}A_{t}$. Let $1\leq t\leq k$. Notice that $\bigcap_{j\in A_{t}}S(Q_{j})$ is $S(P_{t})$-primary by [6, Lemma 4.3]. It is convenient to denote $\bigcap_{j\in A_{t}}Q_{j}$ by $I^{\prime}_{t}$. Thus $(I:s)=S(I)=S(I^{\prime}_{1})\cap S(I^{\prime}_{2})\cap\cdots\cap S(I^{\prime}_{k})$ with $S(I^{\prime}_{t})$ is $S(P_{t})$-primary for each $t\in\{1,\ldots,k\}$ and $S(P_{1}),\ldots,S(P_{k})$ are distinct. After omitting those $S(I^{\prime}_{i})$ such that $S(I^{\prime}_{i})\supseteq\bigcap_{t\in\{1,\ldots,k\}\setminus\{i\}}S(I^{\prime}_{t})$ from the intersection, we can assume without loss of generality that $(I:s)=S(I)=\bigcap\limits_{t=1}^{k}S(I^{\prime}_{t})$ is a minimal primary decomposition of $S(I)$. Next, we claim that $I=(I:s)\cap(I+Rs)$. It is clear that $I\subseteq(I:s)\cap(I+Rs)$. Let $y\in(I:s)\cap(I+Rs)$. Then $ys\in I$ and $y=a+rs$ for some $r\in R$. This implies that $ys=as+rs^{2}$ and so, $rs^{2}\in I$. Hence, $r\in S(I)=(I:s)$. Therefore, $y=a+rs\in I$. This shows that $(I:s)\cap(I+Rs)\subseteq I$. Thus $I=(I:s)\cap(I+Rs)$ and hence, $I=(\bigcap\limits_{t=1}^{k}S(I^{\prime}_{t})\cap(I+Rs))=\bigcap\limits_{t=1}^{k}(S(I^{\prime}_{t})\cap(I+Rs))$. Let $1\leq t\leq k$. For convenience, let us denote $S(I^{\prime}_{t})\cap(I+Rs)$ by $Q^{\prime}_{t}$. As $S(I^{\prime}_{t})$ is $S(P_{t})$-primary with $S(I^{\prime}_{t})\cap S=\emptyset$ and $(I+Rs)\cap S\neq\emptyset$, we obtain from Proposition 7(2) that $Q^{\prime}_{t}$ is $(P_{t})\cap rad(I+Rs)-SS$-primary. Since $S(S(P_{t}))=S(P_{t})$, $S(rad(I+Rs))=R$, it follows that $S(S(P_{t})\cap rad(I+Rs))=S(P_{t})$. Notice that $S(Q^{\prime}_{t})=S(S(I^{\prime}_{t})\cap(I+Rs))=S(I^{\prime}_{t})$, as $S(S(I^{\prime}_{t}))=S(I^{\prime}_{t})$ and $S(I+Rs)=R$. Hence, for all distinct $i,j\in\{1,\ldots,k\}$, $S(S(P_{i})\cap rad(I+Rs))\neq S(S(P_{j})\cap rad(I+Rs))$ and for each $i$ with $1\leq i\leq t$, $S(Q^{\prime}_{i})\nsupseteq\bigcap_{t\in\{1,\ldots,k\}\setminus\{i\}}S(Q^{\prime}_{t})$. Therefore, $I=\bigcap\limits_{t=1}^{k}Q^{\prime}_{t}$ is a minimal $S$-primary decomposition.
Recall from [2], let $R$ be a ring, $S\subseteq R$ a multiplicative set and $I$ an ideal of $R$ disjoint from $S$. Let $P$ be an $S$-prime ideal of $R$ such that $I\subseteq P$. Then $P$ is said to be a minimal $S$-prime ideal over $I$ if $P$ is minimal in the set of the $S$-prime ideals containing $I$. Also Ahmed [2, Remark 2] proves in $S$-Noetherian rings that the set of minimal $S$-prime ideals is finite if $S$ is a finite multiplicative set.
A prime ideal is said to be a minimal prime ideal if it is a minimal prime ideal over the zero ideal. Emmy Noether showed that in a Noetherian ring, there are only finitely many minimal prime ideals over any given ideal [10, Theorem 88]. A natural question arises:
Question 2.1.
Is the collection of minimal prime ideals in $S$-Noetherian rings finite for any multiplicative set $S$?
The answer to the above question is negative. In the following example, we provide an $S$-Noetherian ring that has infinitely many minimal prime ideals.
Example 12.
Consider the ring $R=\dfrac{F[x_{1},x_{2},\ldots,x_{n},\ldots]}{(x_{i}x_{j};i\neq j,i,j\in\mathbb{N})}$, where $F$ is a field. Let $y_{i}={\overline{x}}_{i}$ be the image of $x_{i}$ under the canonical map. Consider the multiplicative set $S=\{y_{1}^{n}:n\in\mathbb{N}\cup\{0\}$} of $R$. Then $(y_{1})\subseteq(y_{1},y_{2})\subseteq\cdots\subseteq(y_{1},y_{2},\ldots,y_{n})\subseteq\cdots$ is an ascending chain of ideals of $R$ which is not stationary. Consequently, $R$ is not a Noetherian ring. Evidently, $R$ is an $S$-Noetherian ring (see [15, Example 4]). Note that if $P\subset R$ is a prime ideal, then there can be at most one $i$ such that $y_{i}\notin P$. It follows that if $P_{i}$ is the ideal generated by all the $y_{j}$ for $j\neq i$, each $P_{i}$ is prime since $R/P_{i}\cong F[x_{i}]$. Clearly, every $P_{i}$ is a minimal prime ideal of $R$. Thus $R$ has infinitely many minimal prime ideals.
Question 2.2.
Under what condition the set of minimal prime ideals is finite in $S$-Noetherian rings?
We give an answer to the above question under a mild condition on the multiplicative set $S\subseteq R$ (see Theorem 14).
Remark 13.
Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $(0)=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary ideal of $R$. If $Z(R/P_{i})\cap\overline{S}_{i}=\emptyset$, where $\overline{S}_{i}=\{s+P_{i}\mid s\in S\}$ and $Z(R/P_{i})$ denotes the set of zero divisors of $R/P_{i}$ for all $i=1,2,\ldots,r$. Then each $P_{i}$ is a prime ideal of $R$. For this, we will show that $(P_{i}:s)=P_{i}$ for all $s\in S$ and $i=1,2,\ldots,r$. It is clear that $P_{i}\subseteq(P_{i}:s)$. Conversely, let $s\in S$ and $x\in(P_{i}:s)$, then $sx\in P_{i}$; so $(s+P_{i})(x+P_{i})=P_{i}$. Thus $x\in P_{i}$ since $Z(R/P_{i})\cap\overline{S}_{i}=\emptyset$. Now, if $P_{i}$ is a $S$-prime ideal of $R$, then by [2, Proposition 1], $(P_{i}:s_{i})$ is a prime ideal of $R$ for some $s_{i}\in S$. Since $P_{i}=(P_{i}:s_{i})$, $P_{i}$ is a prime ideal of $R$.
Theorem 14.
Let $R$ be an $S$-Noetherian ring, $S\subseteq R$ a multiplicative set, and $(0)=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary ideal of $R$. If $Z(R/P_{i})\cap\overline{S}_{i}=\emptyset$, then minimal prime ideals of $R$ are in the set $\{P_{i}\mid i=1,2,\ldots r\}$.
Proof.
Since $(0)=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$, $rad~(0)=P_{1}\cap P_{2}\cap\cdots\cap P_{r}$, where $P_{i}=rad(Q_{i})$ $(1\leq i\leq r)$ is an $S$-prime ideal of $R$ By Remark 13, each $P_{i}$ $(1\leq i\leq r)$ is a prime ideal of $R$. Let $P^{\prime}$ be any minimal prime ideal of $R$. Then $(0)\subseteq P^{\prime}$, and so $rad~(0)\subseteq rad(P^{\prime})=P^{\prime}$. This implies that $P_{1}\cap P_{2}\cap\cdots\cap P_{r}\subseteq P^{\prime}$. By [6, Proposition 1.11(ii)], there exists $j\in\{1,2,\ldots,r\}$ such that $P_{j}\subseteq P^{\prime}$. Since $P^{\prime}$ is a minimal prime ideal, $P^{\prime}=P_{j}$, as desired. ∎
Proposition 15.
@@
On contrary suppose $sQ_{k}\nsubseteq P$ for all $s\in S$ and $k=1,2,\dots,n$. Then for each $k$, there exists $x_{k}\in Q_{k}$ such that $sx_{k}\notin P$ for all $s\in S$. Evidently, $x_{1}x_{2}\ldots x_{n}\in Q_{1}Q_{2}\ldots Q_{n}$. This implies that $x_{1}x_{2}\cdots x_{n}\in Q_{1}Q_{2}\ldots Q_{n}\subseteq\bigcap\limits_{k=1}^{n}Q_{k}\subseteq P$. Since $P$ is an $S$-prime ideal, by [2, Proposition 4], there exist $s^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $s^{\prime}x_{j}\in P$, a contradiction. Hence there exist $s\in S$ and $k\in\{1,2,\ldots,n\}$ such that $sQ_{k}\subseteq P$. Finally, if $P=\bigcap\limits_{k=1}^{n}Q_{k}$, then by above argument there exist $s\in S$ and $k\in\{1,2,\ldots,n\}$ such that $sQ_{k}\subseteq P\subseteq Q_{k}$. ∎
Now we prove $S$-version of the $1st$ uniqueness theorem:
Theorem 16.
$(\textbf{$S$-version of 1st $S$-uniqueness theorem})$. Let $R$ be a ring and let $S$ be a multiplicative set of $R$. Let $I$ be an ideal of $R$ which admits $S$-primary decomposition. Let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary for each $i\in\{1,2,\ldots,n\}$. Then the $S(P_{i})$ are precisely the prime ideals which occur in the set of ideals $S(rad(I:x))$ $(x\in R)$, and hence are independent of the particular $S$-decomposition of $I$.
Proof.
For any $x\in R$, we have $S^{-1}(I:x)=(\bigcap\limits_{i=1}^{n}S^{-1}Q_{i}:_{S^{-1}R}\frac{x}{1})=\bigcap\limits_{i=1}^{n}(S^{-1}Q_{i}:_{S^{-1}R}\frac{x}{1})$, and so $S(I:x)=\bigcap\limits_{i=1}^{n}(S(Q_{i}):x)$. Evidently, $S(rad(I:x))=\bigcap\limits_{i=1}^{n}rad(S(Q_{i}):x)$. Notice that each $S(P_{i})$ is a prime ideal since $S^{-1}P_{i}$ is a prime ideal and contraction of a prime ideal is a prime ideal. Now we show that each $S(Q_{i})$ is $S(P_{i})$-primary. For this, let $x,~y\in R$ such that $xy\in S(Q_{i})$. Then $\dfrac{xy}{1}\in S^{-1}Q_{i}$, $\dfrac{xy}{1}=\dfrac{a}{s}$ for some $a\in Q_{i}$ and $s\in S$. Consequently, there exists $u\in S$ such that $usxy\in Q_{i}$. Since $Q_{i}$ is $P_{i}$-$S$-primary, there exists $t\in S$ such that either $tusx\in Q_{i}$ or $ty\in P_{i}$. It follows that either $\dfrac{x}{1}=\dfrac{tusx}{tus}\in S^{-1}Q_{i}$ or $\dfrac{y}{1}=\dfrac{ty}{t}\in S^{-1}P_{i}$. This implies that either $x\in S(Q_{i})$ or $y\in S(P_{i})$, and rad $S(Q_{i})=S(rad(Q_{i}))=S(P_{i})$. Therefore $S(Q_{i})$ is $S(P_{i})$-primary. Then $S(rad(I:x))=\bigcap\limits_{i=1}^{n}rad(S(Q_{i}):x)=\bigcap\limits_{x\notin S(Q_{i})}S(P_{i})$, by [6, Lemma 4.4]. Suppose $S(rad(I:x))$ is prime; then we have $S(rad(I:x))=S(P_{j})$ for some $j$, by [6, Proposition 1.11]. Hence every prime ideal of the form $S(rad(I:x))$ is one of the $S(P_{j})$. Conversely, for each $i$ there exists $x_{i}\in\left(\bigcap_{j\in\{1,2,\ldots,n\}\setminus\{i\}}S(Q_{j})\right)\setminus S(Q_{i})$ since the $S$-decomposition is minimal. Consequently, for each $i$, we have $S(I:x_{i})=\bigcap\limits_{i=1}^{n}(S(Q_{i}):x_{i})=(S(Q_{i}):x_{i})$ since $(S(Q_{j}):x_{i})=R$ for all $j\neq i$, by [6, Lemma 4.4]. Then $S(rad(I:x_{i}))=S(P_{i})$, by [6, Lemma 4.4]. This completes the proof. ∎
To prove the second uniqueness theorem for the $S$-Noetherian rings, we need the following results:
@@
Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ be an $P$-$S$-primary ideal of $R$. Then $S^{-1}Q$ is $S^{-1}P$-primary and its contraction in $R$ satisfies $tS(Q)\subseteq Q\subseteq S(Q)$ for some $t\in S$.
@@
Clearly, $S^{-1}Q\neq S^{-1}R$ since $Q\cap S=\emptyset$ and $P=rad(Q)$. Let $x\in S(Q)$. Then $\dfrac{x}{1}=\dfrac{a}{s}$ for some $a\in Q$ and $s\in S$. Consequently, there exists $s^{\prime}\in S$ such that $(xs-a)s^{\prime}=0$; whence $ss^{\prime}x=s^{\prime}a\in Q$. As $Q$ is $S$-primary, there exists $t\in S$ such that either $ss^{\prime}t\in rad(Q)=P$ or $tx\in Q$. This implies that $tx\in Q$ since $P\cap S=\emptyset$, and so $tS(Q)\subseteq Q\subseteq S(Q)$. Now since $Q$ is $S$-primary, by [13, Proposition 2.7], $S^{-1}Q$ is a primary ideal of $S^{-1}R$ and $rad(S^{-1}Q)=S^{-1}(rad(Q))=S^{-1}P$. Also since $P$ is an $S$-prime ideal of $R$, by [2, Remark $1$], $S^{-1}P$ is a prime ideal of $S^{-1}R$. Thus $S^{-1}Q$ is an $S^{-1}P$-primary ideal of $S^{-1}R$. ∎
Lemma 18.
Let $R$ be a ring and let $S$ be a multiplicative set of $R$. Let $I$ be an ideal of $R$ such that $I$ admits an $S$-primary decomposition. Let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition where $Q_{i}$ is $P_{i}$-$S$-primary for each $i\in\{1,2,\ldots,n\}$. Let $S^{\prime}$ be a multiplicative set of $R$ such that $S\subseteq S^{\prime}$. If $Q_{1},Q_{2},\ldots,Q_{m}$ are such that $Q_{i}\cap S^{\prime}=\emptyset$ for each $i\in\{1,\ldots,m\}$ and $Q_{i}\cap S^{\prime}\neq\emptyset$ for each $i\in\{m+1,\ldots,n\}$, then $S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$ is a primary decomposition and there exists $t\in S$ such that $tS^{\prime}(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S^{\prime}(I)$.
Proof.
By [6, Proposition 3.11(v)], $S^{\prime-1}I=\bigcap\limits_{i=1}^{n}S^{\prime-1}Q_{i}$. Since $S^{\prime}\cap Q_{i}\neq\emptyset$ for $i=m+1,m+2,\ldots,n$, by [6, Proposition 4.8(i)], $S^{\prime-1}Q_{i}=S^{\prime-1}R$ for $i=m+1,m+2,\ldots,n$. Consequently, $S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$. As $Q_{i}$ is $P_{i}$-$S$-primary and $S\subseteq S^{\prime}$, it follows that $Q_{i}$ is $P_{i}$-$S^{\prime}$-primary for each $i=1,\ldots,m$. Also by Lemma 17, $S^{\prime-1}Q_{i}$ is $S^{\prime-1}P_{i}$-primary for $i=1,2,\ldots,m$. Thus $S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$ is a primary decomposition. Next, let $x\in S^{\prime}(I)$. Then $\dfrac{x}{1}\in S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$. Write $\dfrac{x}{1}=\dfrac{a_{i}}{s_{i}}$ for some $a_{i}\in Q_{i}$ and $s_{i}\in S^{\prime}$. This implies that $(xs_{i}-a_{i})s=0$ for some $s\in S^{\prime}$; whence $s_{i}sx=sa_{i}\in Q_{i}$ for $i=1,2,\ldots,m$. Then there exists $t_{i}\in S$ such that $t_{i}x\in Q_{i}$ since $Q_{i}$ is $P_{i}$-$S$-primary and $P_{i}\cap S^{\prime}=\emptyset$ for all $i=1,2,\ldots,m$. Put $t=t_{1}t_{2}\ldots t_{m}$. Then $tx\in\bigcap\limits_{i=1}^{m}Q_{i}$, and hence $tS^{\prime}(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S^{\prime}(\bigcap\limits_{i=1}^{m}Q_{i})=\bigcap\limits_{i=1}^{m}S^{\prime}(Q_{i})=S^{\prime}(I)$. Hence $tS^{\prime}(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S^{\prime}(I)$. ∎
For an ideal $I$ of $R$, let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary. Then the $S$-prime ideals $P_{i}$ are said to belong to $I$. Also we say that $P_{i}$ ($1\leq i\leq n$) is isolated $S$-prime if $sP_{j}\nsubseteq P_{i}$ for all $s\in S$ and for all $j\neq i$. Other $S$-prime ideals are called embedded $S$-prime ideals.
@@
($S$-version of 2nd uniqueness theorem). Let $I$ be an ideal of $R$ which admits $S$-primary decomposition, that is, $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition of $I$, where $Q_{i}$ is $P_{i}$-$S$-primary. If $\{P_{{1}},\ldots,P_{{m}}\}$ is a set of isolated $S$-prime ideals of $I$ for some $m$ $(1\leq m\leq n)$, then $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition.